LABORATORY AND TENTATIVE ASTRONOMICAL DETECTION OF THIOCYANIC ACID, HSCN

\underline{SANDRA BRÜNKEN}, ZHENHONG YU, MICHAEL C. MCCARTHY, CARL A. GOTTLIEB, PATRICK THADDEUS, Harvard Smithsonian Center for Astrophysics and School of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138; ARNAUD BELLOCHE, KARL M. MENTEN, Max-Planck-Institut für Radioastronomie, 53121 Bonn, Germany.

We present the first measurements of the rotational spectrum of thiocyanic acid (HSCN) in the microwave and mm-wave regions. HSCN is an isomer of the spectroscopically and astronomically well-studied molecule isothiocyanic acid (HNCS). HSCN has been calculated to lie only 4-14 kcal/mol higher in energy than the most stable isomer HNCS, yet to date HSCN had only been characterised experimentally by matrix-IR spectroscopy. We now succeeded to produce HSCN in a discharge of H_{2}S and either (CN)_{2} or CH_{3}CN in sufficient amounts for its spectroscopic investigation in the gas phase. Three a-type rotational transitions in the K_{a} = 0 ladder and four in the K_{a} = 1 ladder were measured in the frequency range 10 − 35 GHz with our Fourier Transform Microwave Spectrometer (FTM) in a molecular beam. The identification was confirmed by the observation of 6 singly and doubly substituted isotopic species at predicted isotopic shifts. Furthermore, we were able to resolve and analyse the distinctive hyperfine structure due to ^{14}N and D in all species containing these nuclei, providing additional evidence for the identification. An experimental structure derived from the isotopic measurements will be presented. For the main isotopic species additional ΔK_{a} = 0 transitions were observed in selected frequency regions up to 350 GHz with a free-space millimeter-wave absorption spectrometer. Lines of HSCN were surprisingly strong through a dc discharge of H_{2}S and (CN)_{2}, allowing to observe spectra up to J = 30 and K_{a} = 7, as shown in the figure. We also report a tentative detection of HSCN towards SgrB2(M) via four K_{a} = 0 a-type transitions observed in the 3-mm band with the IRAM 30m telescope. The observations yield an abundance of HSCN only a factor 2-5 lower than that of the lowest energy isomer HNCS.